# RefineCoder: Iterative Improving of Large Language Models via Adaptive Critique Refinement for Code Generation

### v1

@@ 6 Conclusions
In this paper, we propose the Adaptive Critique Refinement (ACR) method to iteratively refine code LLMs using self-generated code and external critique, which breaks away from the traditional teacher distillation paradigm and improves the intrinsic one-off code generation ability. We design a composite scoring system and a selective critiquing strategy. These two components are centred around LLM-as-a-Judge and LLM-as-a-Critic to evaluate and critique self-generated code. This simulates the process where a student solves a problem independently and then refines it by comparing it with the reference answer. We develop the RefineCoder-DS-6.7B and RefineCoder-QW-7B models and demonstrate the effectiveness of iterative ACR on HumanEval (+), MBPP (+), LiveCodeBench, and BigCodeBench-hard. Further studies reveal the impact of the components in ACR and demonstrate its OOD code generation ability.
@@ 7 Limitations
Appendix A Elo Algorithm
Firstly, we initialize Elo score $E^{i}=1000$ for each code response $y^{i}$. Then, we iteratively update the Elo scores by using the relative scores between any two responses. Taking the response pair $y^{i}$ and $y^{j}$ as an example, we obtain their match results $S^{i}$ and $S^{j}$:
$S^{i}=\frac{\textrm{score}_{judge}^{ij}}{\textrm{score}_{judge}^{ij}+\textrm{score}_{judge}^{ji}},$ (8)
$S^{j}=\frac{\textrm{score}_{judge}^{ji}}{\textrm{score}_{judge}^{ij}+\textrm{score}_{judge}^{ji}}.$ (9)
Then we obtain the expected probabilities of winning $P^{ij}$ and $P^{ji}$ for $y^{i}$ and $y^{j}$:
$P^{ij}=\frac{1}{1+10^{\frac{E^{j}-E^{i}}{400}}},$ (10)
$P^{ji}=1-P^{ij}.$ (11)
Finally, we update the Elo score:
$E^{i}_{new}=E^{i}+K\times(S^{i}-P^{ij}),$ (12)
$E^{j}_{new}=E^{j}+K\times(S^{j}-P^{ji}).$ (13)
where $K=32$ is a adjustment factor. After the iteration, we obtain the Elo score for each code response and then normalize it to derive the final judge score:
$\textrm{score}^{i}_{judge}=\frac{E_{i}-\min(E)}{\max(E)-\min(E)},$ (14)
where $i\in[0,n]$ and $\textrm{score}^{i}_{judge}\in[0,1]$.
Appendix B Prompts for Judge and Critic
The prompt for pairwise LLM-as-a-Judge and LLM-as-a-Critic are shown in Figure 7 and  8.
Appendix C Prompts for constructing SFT Dataset
We called GPT-4o three times to create the SFT dataset, with the following prompts used:
@@ Appendix C Prompts for constructing SFT Dataset
$\circ$ Prompts for evolving existing instructions, as shown in Figures 10,  11,  1213141516,  17,  18.
@@
Appendix D Prompts for Evaluation
For Humaneval (+) and MBPP (+), we use the prompts designed by OpenCodeInterpreter Zheng et al. (2024a); for Livecodebench and Bigcodebench-hard, we use the official prompts. When conducting evaluations with execution feedback, we have designed a prompt as illustrated in Figure 19. When executing human feedback, following OpenCodeInterpreter, we generate two prompts: one for generating improvement suggestions and the other for refine code according to human feedback, as shown in Figures 21 and  20.
Appendix E Data Leakage Analysis
@@
Appendix F Scaling Law of Iterative ACR
We study the performance trends of RefineCoder-DS-6.7B obtained using ACR and the standard SFT method as the dataset size increases. For the SFT method, we directly trained the model using four datasets of different sizes. For RefineCoder series, we employ an iterative training strategy. For example, after obtaining RefineCoder-DS-10k by fing-tuning on the 10k dataset, we apply ACR to update this 10k dataset, and then randomly sample another 10k from the data pool to form a 20k dataset. The model is fine-tuned again to obtain RefineCoder-DS-20k. The experimental results are shown in Figure 6, as the dataset size increases, our model achieves greater performance improvements compared to the SFT model on HumanEval and MBPP. Moreover, the performance gap between the two models does not show any signs of narrowing, underscoring the impressive scaling capabilities of the iterative ACR method.
@@ Appendix F Scaling Law of Iterative ACR
Figure 6: Performance curves of RefineCoder obtained using iterative ACR and the standard SFT model as the dataset size increases.
Figure 7: The prompt template for pairwise LLM-as-a-Judge.
Figure 8: The prompt template for pairwise LLM-as-a-Critic.

### v2

@@ 6 Conclusions
In this paper, we propose the Adaptive Critique Refinement (ACR) method to iteratively refine code LLMs using self-generated code and external critique, which breaks away from the traditional teacher distillation paradigm and improves the intrinsic one-off code generation ability. We design a composite scoring system and a selective critiquing strategy. These two components are centered around LLM-as-a-Judge and LLM-as-a-Critic to evaluate and critique self-generated code. This simulates the process where a student solves a problem independently and then refines it by comparing it with the reference answer. We develop the RefineCoder-DS-6.7B and RefineCoder-QW-7B models and demonstrate the effectiveness of iterative ACR on multiple code generation benchmarks. Further studies reveal the impact of the each component in ACR.
@@ 7 Limitations
Appendix A Prompts for Judge and Critic
The prompt for point/pair-wise LLM-as-a-Judge and LLM-as-a-Critic are shown in Figure 8,  9 and  10.
Appendix B Prompts for Constructing SFT Dataset
We called GPT-4o to create the SFT dataset, with the following prompts used:
@@ Appendix B Prompts for Constructing SFT Dataset
$\circ$ Prompts for evolving existing instructions, as shown in Figures 12,  13,  1415161718,  19,  20.
@@
Appendix C Prompts for Evaluation
For Humaneval (+) and MBPP (+), we use the prompts designed by OpenCodeInterpreter Zheng et al. (2024a); for Livecodebench and Bigcodebench-hard, we use the official prompts. When conducting evaluations with execution feedback, we have designed a prompt as illustrated in Figure 21. When executing human-like feedback, following OpenCodeInterpreter, we generate two prompts: one for generating improvement suggestions and the other for refine code according to human-like feedback, as shown in Figures 22 and  23.
Appendix D Data Leakage Analysis
@@
Appendix E Diminishing Returns
As shown in Figure 5, we present the proportions of two types of data during the iterative process of RefineCoder-QW-7B. The self-generated code struggles more to outperform as iterations continue, indicating diminishing potential for further gains through self-refinement. The corresponding experimental results are shown in Figure 6. As observed, the model exhibits clear diminishing returns after the third iteration. This suggests that the model may have entered a mild overfitting stage, where further refinement not only fails to improve performance but may also exacerbate the issue.
Figure 5: Proportions of two types of data in the iterative process.
Figure 6: The average Pass@1 value of RefineCoder-QW-7B on HumanEval(+), MBPP(+) and BigCodeBench.
@@ Appendix E Diminishing Returns
Figure 8: The prompt template for pointwise judgment.
Figure 9: The prompt template for pairwise judgment.
Figure 10: The prompt template for pairwise critique.
@@
Figure 23: The prompt template for Human-like Feedback.